An investor was considering stock purchase in two companies. He determined

the following payoff table contingent upon whether company A or B wins a competition for

share of the market.

State a1 (Investment in A) a2(Investment in B)

A wins 30 0

B wins 40 80

What is the Maximax and Maximin strategy? At what probability of A’s winning is the two value Strategies?  


1
Expert's answer
2021-02-24T06:35:49-0500

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 State& a_1 & a_2 \\ \hline
 A\ wins & 30 & 0 \\
 \hline
 B \ wins & 40 & 80
\end{array}"

"\text{strategy }a_1\text{(Investment in A) is Maximin}"

"\text{the ability to get a minimum income of 30 for any outcome}"

"\text{strategy }a_2\text{(Investment in B) is Maximax}"

"\text{probability of earning 80 income with a favorable outcome}"

"\text{probability of A’s winning}"

"p*30+40*(1-p)=0*p+80*(1-p)"

"70p=40"

"p=\frac{4}{7}"

Answer:"a_1\text{ is Maximin};a_2\text{ is Maximin};p=\frac{4}{7}"